PRISM logo

Polar Radar for Ice Sheet Measurements

Curved ice line
Jagged line
Photo Pages Data Pages Information Pages

small logo

Data Page from Oct. 6 - Oct. 13, 2000

All information is from the Australian National Antarctic Research Expeditions Voyage 1, 2000/2001 Davis-Mawson-Heard Island taken by Aurora Australis.

All observations taken at 0500 Universal Time. All temperatures are in degrees Celsius.


Air Temp

Sea Temp
Ice Cndtns
Weather Cndtns
Sea Cndtns
Oct. 6
-46 50.16
142 11.38
partly cloudy
moderate seas
Oct. 7
-46 53.2
142 02.1
None partly cloudy very rough seas with heavy swell
Oct. 8
-50 02.05
141 53.94
None cloudy, fine and clear rough seas with heavy swell
Oct. 9
-53 37.0
141 48.8
None overcast rough seas with heavy swell
Oct. 10
-53 26.80
140 39.5
None overcast with passing snow showers very rough seas with heavy swell, vessel pounding heavily.
Oct. 11
-53 44.735
141 45.395
None fine and clear rough seas with heavy swell
Oct. 12
-53 59.78
140 31.03
None overcast, fine and clear moderate to rough sea with a moderate to heavy confused swell
Oct. 13
-54 44.9
133 59.3
None --- moderate to rough seas with a moderate swell


Oct 6: Expos put in a magnificent effort until last light looking for sediment trap. Conducted search pattern overnight based on strength of radio signal from unit. Visual and radio search continuing -'good buoys are hard to find!"

Oct. 7: Buoy oh Buoy -we got it!! Sharp-eyed Judy Clarke spotted the sediment trap assembly buoys at 1730 (local Friday 6 October) -by 1900 the acoustic release and floats were onboard, followed by a current meter and sediment trap at 2000, another current meter and sediment trap at 2116 and the final ones at 2200. All were prepared for redeployment overnight.

This commenced at 1330 (Sat 7 Oct local) at 46 degree 47.91'S; 142 degrees 06.87'E and should be completed by 1630. Info recorded in Event Log. We are very pleased to retrieve this important data and equipment. Many thanks to the dedicated band of lookouts who assisted in the operation and to the ships' crew for their efforts in bringing it all onboard.

Field Training commenced today. Other programs continue. Also preparing for deployment of Continuous Plankton Recorder after completion of sediment trap work.

Oct 8: Completed deployment of sediment traps (2 out of 3 put out due to worsening weather) at 1615 (local Saturday 7 October). Continuous Plankton Recorder in the water by 1630. Ship based Home Page established (thanks to Andrew Klek. and Dave Rasch), which is most sophisticated with features such as track charts that automatically update every hour, safety info, sitreps, and photos,�

Safety Muster held this morning followed by fire safety talk and demonstration of use of the stretcher. Field Training continues with rope and knot work on the agenda.

Oct. 9: Heading towards sediment trap release site. Expect to arrive later today. Currently preparing to bring Continuous Plankton Recorder back onboard.

Standing room only for talk on the LIDAR by Andrew Klekociouk. Onboard programs continue.�

Monday's Fuel Figures: 1,153.633�

Oct. 10: Winds have eased from 55 knots to 35. Remain in vicinity of sediment trap in hope of better weather tomorrow so can retrieve them. Onboard programs continue. Field training postponed due to conditions which have resulted in a quiet ship yet again.

Oct. 11: Remained in vicinity of sediment trap overnight in anticipation of good weather window today. Winds dropped and seas calmed so sediment trap released at 1510 (local). Radio transponder must not be working so are continuing visual search in grid pattern. All hands on deck looking.

Oct. 12: The sediment traps didn't have a chance of escaping us!! Yesterday afternoon we stationed all 80 + expos at vantage points around the ship. Keen eyed marine geologist Jason Whitehead spotted the buoys at� 1710, by 1730 the transmitter was onboard, the first sediment trap at 1750, the second at 1845, and retrieval complete at 1915. The Marine Science team prepared the traps and wires overnight for redeployment this morning. Redeployment started at 0810 and was completed at 0947 with the traps located at 53 44.549 degrees South; 141 45.685 degrees East. The Continuous Plankton Recorder then commenced its second transect. Now heading for Davis Ice Edge. Clocks back one hour tonight. Thanks to everyone for their patience and good humour.�

Oct. 13: Field Training has resumed. Initial Planning Meeting for Helicopter Operations held. Ship quiet as expeditioners are working on costumes for Fright Night" for Friday the 13th. No black cats seen yet and we are avoiding ladders.




Photo Pages Data Pages Information Pages



PRISM © 2002, 2003 is brought to you by
NSF logo
National Science Foundation University of Kansas
NASA logo
KTEC logo
Kansas Technology Enterprise Corporation
University of Kansas logo
University of Kansas